Jack wants to make 30 quarts of a 15% solution of weed killer for his lawn. He has some 12% solution and some 20% solution. How much of each does he need?
Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6